OpenStudy (anonymous):

What is the sum of 3.6 × 103 and 6.1 × 103?

OpenStudy (anonymous):

370.103 is the first one

OpenStudy (anonymous):

and the second one is 628.3

OpenStudy (anonymous):

\[3.6\times 103+6.1\times 103=9.7\times 103\] and a calculator

OpenStudy (anonymous):

9.7 × 10^3 9.7 × 10^6 97 × 10^3 97 × 10^6

OpenStudy (anonymous):

@satellite73 @strawberry12345

OpenStudy (anonymous):

@amistre64 @AnswerMyQuestions @amistre64 @ayyookyndall @Anon101